Tue 30 Nov - All Systems Ran Without Problems
Mon 29 Nov - Microsoft SharePoint Upgrade & Outgoing Email No Longer Blocked
Microsoft SharePoint Upgrade
The university is beginning a programme to upgrade the existing Microsoft SharePoint websites to SharePoint 2010. The first stage of this upgrade will commence this evening at 10pm and take approximately three hours. During this time the SharePoint servers will be unavailable and a "server under "will be displayed to users attempting to access the servers.
Further upgrade work will be carried out on Tuesday and Thursday evenings this week.
More information is available here.
Outgoing Email No Longer Blocked
Last week Lancaster emails were being blocked from other email service providers, such as Microsoft. These problems have been fixed and all Lancaster email is now being delivered to Google/Gmail, AOL and Microsoft Live/Hotmail/MSN. If any of the mail you send is returned marked as undeliverable please contact the ISS Service Desk.
For more information on how to recognise and avoid phishing scams please follow the link: more information
Sat 27 Nov - Resilience Testing
To ensure ISS is able to provide a secure datacentre in the event of an unexpected power loss all back-up infrastructure is tested regularly. Thus, in the event of power being lost all essential ISS data centre equipment remains in a powered up state and users experience no connectivity problems or data loss. The data storage back-up equipment was tested on 27th November in the ISS Data Centre
Users across campus should not be affected.
Fri 26 Nov - Ongoing compromised account problems
Thu 25 Nov - compromised account problems
The University continued to experience problems directly related to the compromising of IT accounts.
Wed 24 Nov - Ongoing Compromised Account problems
Email from the University was today blocked by a very popular global provider of firewalls, Barracuda Networks. The University has since sent requests to Barracuda to be removed from the list of known sources of Spam and this was accepted at 12 noon. Unfortunately, unblocking can take between 12 - 24 hours to propagate the Barracuda network. A slightly more technical explanantion follows.
The Lancaster University IP address has been seen by Barracuda Networks to transmit spam email in the past 30 days and consequently we have been given a "poor" rating. This has lead directly to the blocking of University email. The University IP address will be automatically rechecked several times each day and may be adjusted again if other issues are observed.
The Barracuda Spam Firewall applies rules to all email sent from an IP address known to Barracuda Networks to have a "poor" rating. The Firewall has an option to decline email from these IPs. This is an option that the Administrator of the Barracuda Spam Firewall may enable and if it is enabled, email from the IP address is blocked. That is what has happened to Lancaster University.
Removal of the blocking takes between 12 and 24 hours for changes to globally propagate to all Barracuda Spam Firewalls.
Outgoing email sent to Google and AOL is being delayed and totally blocked by Tiscali. This is due to another user disclosing their password. Email sent to Microsoft accounts is currently being received without problems.
Tue 23 Nov - Email Remains Blocked, all other systems ran without problems
Software upgrades and network maintenance
During the Network At Risk period new software was installed on wireless system components. No wireless network services were available whilst the maintenance took place.
Connections to Chancellors Wharf and Cable Street residences were unavailable for brief periods whilst work to improve resilience was carried out. Work began at 0700 and completed by 0800.
There were short interruptions to network services between 0700 and 0830 in the Engineering, Whewell, George Fox & Pendle buildings whilst network maintenance work was undertaken.
All maintenance works were completed successfully.
Mon 22 Nov - Blocked Email
Unfortunately, and despite numerous advisory warnings, a number of people haved revealed their username and password to recent phishing attackers and their accounts have subsequently being used to send out Spam. As a direct consequnece, Microsoft blocked email from Lancaster University. More specifically, outgoing University email was not received by Microsoft Hotmail, Microsoft Live and off-site Microsoft Exchange email addresses. The ISS mail team was, and is working to get the University unblocked.
More information is available here.
Sat 20 Nov - Resilience Testing
Resilience testing in the Faraday datacentre was completed successfully. Full service was maintained throughout the period of testing.
Fri 19 Nov - Recent Phishing Attack Hijacks University Email
Some people have responded to the recent phishing attack, subject: "Your mailbox has exceeded the allowed storage". Their accounts are being used to send out Spam one by one. If you think that you may have responded to this attack please contact the ISS Service Desk immediately. More Information
Thu 18 Nov - Fake Webmail Messages
Fake email messages subject: "Your mailbox has exceeded the allowed storage" are arring on campus. Do NOT follow the link in the message. More Information
Wed 17 Nov - LUVLE Maintenance
Maintenance will be performed on Wednesday to apply the latest software updates to LUVLE.
The LUVLE Spaces that will be affected by the maintenance period include AC.F, ECON, PSYC and LEC.
Tue 16 Nov - Network Maintenance in County South and Bowland Tower
Network maintenance will take place in County South between 0730 and 0830. Some network resources will be unavailable for a short period (under 20 minutes) while new equipment is connected. Additionally, network maintenance in the Bowland Tower area during this time will cause the temporary loss of all network resources for approximately 10 minutes. This will affect Slaidburn House, parts of the School of Health and Medicine and B and C floor in Alexandra Square.
Mon 15 Nov - LUVLE Software Update Applied
The latest software update was tested and applied between 6pm and 6.30pm on Saturday. Unfortunately the update (Hotfix4) seriously affected service on one of the four LUVLE servers. Full service was restored this morning. More Information
Sat 13 Nov - LUVLE Update
Another software update from IBM (hotfix 4) was applied to improve the availability of LUVLE . The update was implemened between 6pm and 6.30pm on Saturday.
Fri 12 Nov - All systems ran without problems
Thu 11 Nov - All systems ran without problems
LUVLE - Update: LUVLE now appears to be stable. ISS and IBM are continuing with their investigation to establish the root cause of the problems experienced recently.
Wed 10 Nov - All systems ran without problems
Tue 09 Nov - All systems ran without problems
Mon 08 Nov - All systems ran without problems
Fri 05 - Network Outage - Bowland Tower Area
A network outage occured in the Bowland Tower area and affected Slaidburn House, parts of the School of Health and Medicine and B and C floor in Alexandra Square.
Thu 04 - AV Problem in Lecture Theatres RESOLVED
The AV equipment in the lecture threatres is now functioning normally.
Wed 03 - LUVLE Logon Problem and Issue With AV Equipment in Lecture Theatres
Due to an unannounced AMX programming update rolled out by the supplier (ProAV) yesterday afternoon, there is an ussue with AV equipment in:
- LUMS 4, 5, 6, 7, 9, 10, 11, & 12
- Furness LT 1, 2, & 3
- Fylde LT 1, 2, & 3
A member of the ISS AV staff will be on hand to ensure that the equipment CAN be used for lectures. If you have a problem with the AV equipment ring the ISS Service Desk on 01524 5 10987. The supplier is onsite attemping to rectify the problem.
Updates were applied during the night to fix a logon problem which was affecting LUVLE Spaces in Ac.F, LEC, Economics and Psychology.
Tue 02 Nov - LUVLE Update
IBM have been working with ISS to solve the recent LUVLE hardware problems and have now applied a software update on the LUVLE hardware to make it more stable.
Follow link for more information.
Mon 01 Nov - LUVLE Scheduled Maintenance
The LUVLE service will be interrupted for up to an hour on Monday 1st November, from 17:30 to 18:30, to allow for the installation of the anticipated Hotfix3 from IBM. We will notify you if this period is delayed or cancelled.